OpenStudy (anonymous):

Use matrix Method to prove the following lines are concurrent 2x+5y+19 = 0 4x-3y-1=0 5x-2y+4=0

OpenStudy (anonymous):

find the determinant. It should be =0 if this lines are concurrent

OpenStudy (anonymous):

? can you be a little more specific please ?

OpenStudy (anonymous):

Determinant of the coefficient matrix.

OpenStudy (anonymous):

... little bit more please.

OpenStudy (anonymous):

If a matrix is singular it is not invertible.

OpenStudy (anonymous):

? little bit more please :)

OpenStudy (anonymous):

It is helpful if you let us know what your thoughts on it are so far; this will prevent people from telling you things you already know, and help them write their answers at an appropriate level-Thanks.

OpenStudy (anonymous):

im not to sure. but is the determinant like the 2 and 5 ?

OpenStudy (anonymous):

you look preaty raw on this topic, just read something about matrix algebra and it's aplication to linear system of equations

OpenStudy (anonymous):

I am not sure either of what you are trying to say, there is only one possible determinant.

OpenStudy (mani_jha):

Like FFM said, write down the coefficients of x, y and z in separate columns of a determinant(3 rows and 3 columns). Then find the determinant's value(by expansion). If it comes out to be non-zero, the lines are concurrent.

OpenStudy (anonymous):

So the co-efficents are 2, 5, 19 4,3, 1 5, 2 ,4 ?

OpenStudy (mani_jha):

2,5,19 4,-3,-1 5,-2,4 This would be the determinant. Can you evaluate it by expansion?

OpenStudy (anonymous):

"If it comes out to be *zero*, the lines are concurrent."

OpenStudy (mani_jha):

But if determinant is zero,then the matrix is not invertible, so it wouldn't be possible to solve it!

OpenStudy (anonymous):

LOL, so if its zero, they are all concurrent, thats it ?

OpenStudy (anonymous):

We don't need to solve it. If it's non-zero then it will give the area of the corresponding parallelogram.

OpenStudy (anonymous):

Here, the rank is 2, this matrix is going to have non-trivial solutions.

OpenStudy (anonymous):

think about it this way. Since this determinant is giving the are of a triangle formed by this lines, if the lines are concurrent they form no triangle, so area is 0

OpenStudy (zarkon):

I would row reduce the matrix.

OpenStudy (mani_jha):

Oh you're right. Thanks FFM and Myko!

OpenStudy (zarkon):

a zero determinant does not imply concurrent

OpenStudy (zarkon):

it is possible that there is no solution (though this one has a solution) no solution could give a zero determinant

OpenStudy (mani_jha):

@Zarkon, In this case the determinant does come out to be zero. I was thinking that if the lines are concurrent, it means that they have a unique solution. So determinant would be non-zero.

OpenStudy (zarkon):

there is a unique solution...you have an augmented 3x3...not a system of 3 equations and 3 unknowns

OpenStudy (zarkon):

the determinant should be be used in this type of problem...one should use elementary row operations to get the matrix in rref

OpenStudy (zarkon):

*should NOT be used

OpenStudy (anonymous):

sry, but i desagree. Determinant makes it much easyer and faster. :) YTou can find a proof of this anywhere. Look here for example: http://www.math.wisc.edu/~robbin/461dir/coordinateGeometry.pdf

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@Mani_Jha

OpenStudy (anonymous):

so. It would be 2 5 19 2 -3 -1 5 -2 4 then what do i multiply it with ? im still confused.

OpenStudy (anonymous):

just find the determinant. If it's 0 lines are concurrent, if not--- not

OpenStudy (anonymous):

i got a determinant of -506 ? did i do something wrong lol ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

SEC I CHECK

OpenStudy (anonymous):

you are wrong it's equal 0 |2 5 19| |4 -3 -1| = -24-152 -25 +285 -80 -4 = 0 |5 -2 4|

OpenStudy (anonymous):

@axle1

OpenStudy (anonymous):

awesome, thanks for the help man! really appreciate it :)

OpenStudy (anonymous):

The OP probably meant collinear. This system clearly has non-trivial solutions.

OpenStudy (zarkon):

@myko suppose that you have 3 equation similar to the above... it is possible that once you put it in a matrix like you did here you could end up with \[\left[\begin{matrix}1 & 1 & 0\\ 0 &0 & 1\\0 & 0 & 0\end{matrix}\right]\] for example the system below x+y=1 2x+2y=3 x-y=1 this system has no solution and thus is not concurrent but its determinant is zero

OpenStudy (zarkon):

I ment the system x+y=1 2x+2y=3 x+y=2

OpenStudy (anonymous):

it's a special case, becouse your first 2 columns are dependent, so span same line

OpenStudy (anonymous):

But the concurrent means unique solution right?

OpenStudy (anonymous):

right

OpenStudy (zarkon):

yes..unique solution

OpenStudy (zarkon):

the above does not have a unique solution...thus a determinant of zero does not tell me that are concurrent

OpenStudy (anonymous):

Here we don't have non-trivial solutions so this can't be concurrent ..

OpenStudy (anonymous):

I mean we have.

OpenStudy (zarkon):

and when I say the above...I mean the equatiuon i made up

OpenStudy (anonymous):

I was referring to the OP's system.

OpenStudy (zarkon):

the original equation of the op has a unique solution...but not because the determinant is zero

OpenStudy (anonymous):

x+y=1 2x+2y=3 x+y=2 you don't get that from the matrix you wrote....

OpenStudy (zarkon):

rew reduce \[\left[\begin{matrix}1 & 1 & 1\\ 2 &2 & 3\\1 & 1 & 2\end{matrix}\right]\]

OpenStudy (zarkon):

that is the augmented matrix

OpenStudy (anonymous):

I don't understand, how come the system has unique solution? This is homogeneous system, for unique solution it implies the trivial solution. But the rank of the coefficient matrix is 2 implying non-trivial solutions, ergo making non-unique solution. where I am going wrong?

OpenStudy (zarkon):

i need to learn to spell...*row reduce

OpenStudy (zarkon):

row reduce the original matrix...you will get a unique solution

OpenStudy (zarkon):

x=-2 and y=-3

OpenStudy (anonymous):

Zarkon, for a homogeneous system, if the rank of the augmented matrix is less than the number of unknowns that implies non-trivial solution. By Rouché–Capelli theorem.

OpenStudy (zarkon):

you don't have a homogenous system

OpenStudy (anonymous):

Crap! Sorry I thought this system has 3 unknowns :(

OpenStudy (zarkon):

are we done now :)